OpenStudy (anonymous):

With a tailwind, a jet flew 2000 miles in 4 hours. The jet's return trip against the same wind required 5 hours. Find the jet's speed and the wind speed.

OpenStudy (anonymous):

(jetspeed + windspeed) * 4 = 2000 = (jetspeed - windspeed)*5

OpenStudy (jamesj):

Suppose that with no wind the jet's speed is v. Let the wind have speed w. For the first leg of the trip, the speed of the jet is v + w because the wind is assisting the jet. Now, speed = (distance)/(time) hence v + w = 2000/4 = 500 For the return trip, the speed of the jet is v - w because the wind is working against the jet. Hence v - w = 2000/5 = 400 Therefore you have two simultaneous equations: v + w = 500 v - w = 400 Now solve them to find v and w.